# The Energy Barrier of Precritical Nuclei in Hard Spheres is Consistent with Predictions

### v3

@@ Acknowledgements.
It has been a pleasure to discuss the challenge of the hard sphere nucleation problem with Marjolein Dijkstra, Daan Frenkel, Peter Harrowell, John Russo and Thomas Speck. The authors gratefully acknowledge the Agence National de Recherche grant DiViNew for support. This work benefited from the technical contribution of the joint service unit CNRS UAR 3750.References
@@ Appendix B Barriers
The outcome of the bond-orientational order parameter analysis is a discrete value of the free energy barrier for a certain cluster size (Fig. 3). These discrete values were binned manually to ensure a reasonable representation of the raw data. The size of the bins varies according to the number of clusters found for each size of nuclei. The error bar is then calculated as the square root of the number of clusters in each bin.

### v4

@@ Acknowledgements.
It has been a pleasure to discuss the challenge of the hard sphere nucleation problem with Marjolein Dijkstra, Daan Frenkel, Peter Harrowell, Laura Filion, John Russo and Thomas Speck. The authors gratefully acknowledge the Agence National de Recherche grant DiViNew for support. This work benefited from the technical contribution of the joint service unit CNRS UAR 3750.References
@@
- Wöhler and Schilling [2022] W. Wöhler and T. Schilling, Phys. Rev. Lett. 128, 238001 (2022).
@@
- Russo et al. [2013] J. Russo, A. C. Maggs, D. Bonn, and H. Tanaka, Soft Matter 9, 7369 (2013).
- Wood et al. [2018] N. Wood, J. Russo, F. Turci, and C. P. Royall, J. Chem. Phys. 149, 204506 (2018).
@@ Appendix B Barriers
The outcome of the bond-orientational order parameter analysis is a discrete value of the free energy barrier for a certain nucleus size. These discrete values were binned manually to ensure a reasonable representation of the raw data (Fig. 6). The size of the bins varies according to the number of nuclei found for each size. The error bar is then calculated as the square root of the number of nuclei in each bin.